OpenStudy (anonymous):

multiply and simplify if possible sqrt of 6 times sqrt2

OpenStudy (anonymous):

So do you have a calc?

OpenStudy (anonymous):

\[\sqrt{6} x \sqrt{2}\]

OpenStudy (anonymous):

not close to me

OpenStudy (anonymous):

Ok hold on while I find a calculator that will help you

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

To be correct is your equation 1.\[\sqrt{6}x \sqrt{2}\]? Or is it 2\[\sqrt{6}\times \sqrt{2}\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok Thank you.

OpenStudy (anonymous):

This might help you. It is the question answered in the ways you might need it. http://www.mathway.com/problem/NTAyNDIxNjAz

OpenStudy (anonymous):

thank you heres a medal

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

And please close this topic. @jeannacrisp